Shift Summary for Eve shift, 06/08/99 (fwd)

Ashkan_Alavi-Harati (alavi@fnal.gov)
Wed, 9 Jun 1999 00:36:23 -0500

Shift Summary for Eve shift, 06/08/99

On Shift: Valmiki Prasad Ashkan Alavi

Hours of beam delivered at all intensities: 1.0
Hours of beam delivered at nominal intensity: 0
Hours of E799 data (special runs not included): 0
Hours of special runs(laser scans, muon runs...): 3.0

Accounting of lost time: Various problems with the accelerator.

========================================================================
Runs ended this shift:

13443: KTEV_PED
13444: Laser scan
13450: Cosmics
13451: Cosmics
13454: BEAM_MUON_4E11 (standard setting)
13455: BEAM_MUON_4E11 (Magnet on at 100 MeV/c PtKick)

Tapes brought to FCC this shift: 10
QKC001-0003, QKS001-003, QKK001-003, QKB001.

========================================================================
Accesses during shift:

Elizabeth and Sandrine completed their access from the last shift.

Rick K. and Ashkan to diagnose the SDAQ ADC on CAMAC crate 23.

Rick K. and Ashkan removed the bad bypass module from CAMAC crate 23
in station 4.

========================================================================
Description of Shift:

15:30 Still waiting for beam. No new message on the screen.

15:40 ET and SL adjusted CA dc-offs on a controlled access.

15:45 Ed and the present crew in the control room discussed the shift plan
which is taking muon runs with the analyzing magnet off for dc alignment,
and with the magnet on for chamber studies.

16:30 A pedestal run taken to update MA and CA peds.

18:23 Slow DA shows error in the read out of the ADC in CAMAC crate 23.

18:30 DN installed two new Ar-Ethan bottles on West side.

19:30 JH moves the HCC-RF timing 1ns early. VP adjusts HV of channel
2362. New CsI electron calibration constants are installed as well.

20:30 The problem in crate 23 is tracked down to the bad bypass module
which is bad and as a result DA can not communicate with the crate.

21:30 RK and AA removed the bad bypass module on a controlled access. The
controller in crate 23 is fine.

22:50 Beam returns. Setting up for a straight thorough Muon run.

23:30 Old problem, can not append to tapes after a dart_fresh-start. New
tapes were mounted in the drives.

23:35 Muon run 13454 with magnet off starts.
23:43 run ends.

23:53: Start a special Muon run with magnet on (ptkick=100 Mev/c).

-------------------------------